Chapter 10

What is the THR for a 45-year-old at 75% of maximum? Show work. (3 points)


What are the protein needs for 150 lb. sedentary female? Show work. (3 points)


Based on a 2000 kcal diet. What is the amount of protein in grams at 15% of 2000 kcal? Show work. (3 pts)
Answer #1

The targeted heart rate is calculated using the Heart Rate Reserve (HRR) method

Here, Age of person = 45 years; 75% maximum

Maximum Heart Rate (MHR) = 220 - 45 = 175 BPM

In 45 year old person, average Resting Heart Rate (RHR) is 80BPM,

HRR = MHR - RHR

HRR = 175 - 80 = 95

Now, To calculate Target heart rate use following formula

THR = [HRR X 0.75 (75%)] + RHR

THR = [95 X 0.75] + 80

THR = 71.25 + 80 = 151.25 BPM

The Target Heart Rate of 45 year old is 151.25 when your desired effort is maximum 75%.